# StructuredField: Unifying Structured Geometry and Radiance Field

### v1

@@ 4.3 Applications
For many applications such as physical simulation and deformation, the tetrahedral mesh representation in our StructuredField has inherent advantages. Unlike triangle mesh, which only represent the surface of an object, tetrahedral mesh have internal structure, making it more suitable for physically accurate simulations. On the other hand, thanks to our carefully designed reparameterization strategy, the primitives in our representation are always confined within the corresponding tetrahedra. This ensures that our representation accurately reflects the mesh deformation. In contrast, previous methods using triangle mesh may suffer from artifacts during deformation, as the 3D Gaussians could shift away from the triangular surfaces, leading to potential inconsistencies between 3D Gaussians and triangle mesh.
@@
After constructing the tetrahedral mesh, we treat it as a mass-spring system, where each vertex is modeled as a mass point, and the edges serve as springs connecting these points. To simulate the physical behavior of the system, we implement the XPBD [37] (Extended Position Based Dynamics) algorithm using Taichi [20] programming language. Specifically, we apply constraints to preserve the spring lengths and tetrahedral volumes during the simulation, ensuring realistic deformations and interactions. Fig. 9 shows the results of our physics simulation: please refer to the supplemental video for dynamic motion. By adjusting various parameters of the object, such as mass and the strength of spring constraints, we can simulate different effects on the same object.
@@
We utilize lattice deformation as a mechanism to drive the deformation of the tetrahedral mesh. As the mesh vertices undergo deformation, we reparameterize the associated Gaussian primitives based on the updated positions of the mesh vertices. As demonstrated in Fig. 10 even under large-scale deformations, our representation still produces reasonable rendering results.
5 Conclusion
We introduced StructuredField, a novel structured 3D representation that unifies high-fidelity rendering and structured geometry. Our main contribution is using reparameterization to make tetrahedral mesh rendering differentiable. An orientation-preserving homeomorphism has been proposed to ensure the mesh quality during optimization. Extensive experiments have been conducted to verify the effectiveness of StructuredField on the aspects of rendering quality, physics simulations and deformation modeling.
@@ 5 Conclusion
Although our proposed StructuredField representation can simultaneously recover high-quality rendering and structured geometry from multi-view images, there are still some limitations: First, during the training process, vertices need to pass through the neural network, which leads to longer training time than unstructured radiance field representation. In addition, since we need to maintain structured geometry during the reconstruction process, our solution space is more restricted than unstructured radiance field representation, so more number of primitives are required to represent the scene to achieve similar rendering effects.

### v2

@@ 4.3. Applications
For many applications such as physical simulation and deformation, the tetrahedral mesh representation in our StructuredField has inherent advantages. Unlike triangle meshes, which only represent the surface of an object, tetrahedral meshes have an internal structure, making it more suitable for physically accurate simulations. On the other hand, thanks to our carefully designed reparameterization strategy, the primitives in our representation are always confined within the corresponding tetrahedra. This ensures that our representation accurately reflects the mesh deformation. In contrast, previous methods using triangle meshes may suffer from artifacts during deformation, as the 3D Gaussians could shift away from the triangular surfaces, leading to potential inconsistencies between 3D Gaussians and triangle meshes.
Figure 10. Gallery of physical simulation results. The original shape is shown in the leftmost column, with the corresponding physical simulations displayed in the right column. Videos are provided in the supplementary materials for further visualization.
@@
After constructing the base tetrahedral mesh $\mathcal{M}$, we use it for physical simulation. This process yields a sequence of vertex positions defining its dynamic state over time, denoted as $\mathcal{M}^{k}$. The motion of any implicitly defined child tetrahedra is then consistently derived from their parent elements within $\mathcal{M}^{k}$. This is achieved by applying the previously optimized barycentric coordinates of the subdivision control points to the vertices of the dynamically deformed parent tetrahedra. This process yields the dynamic rendering mesh, $\mathcal{M}_{r}^{k}$.
To simulate elastic deformations specifically, we treat the base mesh $\mathcal{M}$ as a mass-spring system. In this system, each vertex is modeled as a mass point, and the edges of the tetrahedra serve as springs connecting these points. We implement the XPBD algorithm (Macklin et al., 2016) using the Taichi programming language (Hu et al., 2019). Other types of simulation effects or complementary demonstrations presented are realized in Houdini (SideFX Software Inc., 2023). Fig. 10 shows results from our physics simulation approach; please refer to the supplementary video for dynamic motion.
@@
We utilize lattice deformation as a mechanism to drive the deformation of the tetrahedral mesh. As the mesh vertices undergo deformation, we reparameterize the associated Gaussian primitives based on the updated positions of the mesh vertices. As demonstrated in Fig. 9, even under large-scale deformations, our representation still produces reasonable rendering results.
Level-of-Detail
The hierarchical nature of our implicit subdivision naturally facilitates an efficient Level of Detail (LOD) rendering strategy. This allows the complexity of the rendered scene to be dynamically adjusted, balancing visual fidelity with computational performance. Starting from the rendering mesh $\mathcal{M}_{r}$ , coarser LODs are generated by progressively replacing sets of four child tetrahedra with their respective parent tetrahedron. For attributes defined per tetrahedron, the opacity of the parent tetrahedron is taken as the maximum of its four children’s opacities, and its rotation quaternion is determined by averaging the rotation quaternions of its children. Fig. 11 demonstrates the effects of our LOD approach.
Figure 11. Demonstration of our method’s LoD capabilities on Shelly Dataset and NeRF Synthetic Dataset.
Table 5. Quality metrics of initial mesh and optimized mesh on NeRFSynthetic Dataset.
lego mic ship chair
AR↑ ARG↑ AVR↓ AR↑ ARG↑ AVR↓ AR↑ ARG↑ AVR↓ AR↑ ARG↑ AVR↓
Initial Mesh $0.853\pm 0.086$ $0.829\pm 0.091$ $1.234\pm 0.228$ $0.826\pm 0.105$ $0.800\pm 0.109$ $1.274\pm 0.297$ $0.823\pm 0.118$ $0.798\pm 0.121$ $1.306\pm 0.358$ $0.848\pm 0.097$ $0.824\pm 0.101$ $1.255\pm 0.285$
Ours Mesh $\textbf{0.855}\pm\textbf{0.053}$ $\textbf{0.829}\pm\textbf{0.056}$ $\textbf{1.229}\pm\textbf{0.202}$ $\textbf{0.849}\pm\textbf{0.051}$ $\textbf{0.821}\pm\textbf{0.054}$ $\textbf{1.236}\pm\textbf{0.208}$ $\textbf{0.849}\pm\textbf{0.054}$ $\textbf{0.822}\pm\textbf{0.056}$ $\textbf{1.254}\pm\textbf{0.226}$ $\textbf{0.854}\pm\textbf{0.053}$ $\textbf{0.827}\pm\textbf{0.056}$ $\textbf{1.238}\pm\textbf{0.213}$
Figure 12. Gallery of mesh results on Shelly dataset and NeRF Synthetic dataset.
5. Conclusion
We introduced StructuredField, a novel structured 3D representation that unifies high-fidelity rendering and structured geometry. Our main contribution is using reparameterization to make tetrahedral mesh rendering differentiable. An orientation-preserving homeomorphism has been proposed to ensure the mesh quality during optimization. Extensive experiments have been conducted to verify the effectiveness of StructuredField in the aspects of rendering quality, physics simulations, and deformation modeling.
@@ 5. Conclusion
Although our proposed StructuredField representation can simultaneously recover high-quality rendering and structured geometry from multi-view images, there are still some limitations: First, during the training process, vertices need to pass through the neural network, which leads to a longer training time than unstructured radiance field representations. In addition, since we need to maintain structured geometry during the reconstruction process, our solution space is more restricted than unstructured radiance field representations, so a larger number of primitives are required to represent the scene to achieve similar rendering effects.
Acknowledgements.
This research was supported by the National Natural Science Foundation of China (No.62122071, No.62272433), the Youth Innovation Promotion Association CAS (No. 2018495) and the Fundamental Research Funds for the Central Universities (No. WK3470000021).